# https://neu-vi.github.io/HOI-Diff/

### v2

@@ Abstract
We address the problem of generating realistic 3D human-object interactions (HOIs) driven by textual prompts. To this end, we take a modular design and decompose the complex task into simpler sub-tasks. We first develop a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioned on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches. We also develop an affordance prediction diffusion model (APDM) to predict the contacting area between the human and object during the interactions driven by the textual prompt. The APDM is independent of the results by the HOI-DM and thus can correct potential errors by the latter. Moreover, it stochastically generates the contacting points to diversify the generated motions. Finally, we incorporate the estimated contacting points into the classifier-guidance to achieve accurate and close contact between humans and objects. To train and evaluate our approach, we annotate BEHAVE dataset with text descriptions. Experimental results on BEHAVE and OMOMO demonstrate that our approach produces realistic HOIs with various interactions and different types of objects.
Figure 1: HOI-Diff generates realistic motions for 3D human-object interactions given a text prompt and object geometry. Please see the sup. mat. for video results. Darker color indicates later frames in the sequence. Best viewed in color.
@@ 1 Introduction
The generation of natural and physically plausible 3D HOIs involves humans interacting with dynamic objects in various ways according to the text prompts, thereby posing several challenges. First, the variability of object shapes makes it particularly challenging to generate semantically meaningful contact between the human and object to avoid floating objects. Second, the generated HOIs should be faithful to the input text prompts as there are many plausible interactions between human and the same object (e.g., a person carries a chair, sits on a chair, pushes or pulls a chair). Text-driven 3D HOI synthesis with a diverse set of interactions is still under-explored. Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either HOIs [11, 34] or textual descriptions [3].
@@
On one hand, recent methods [23, 17, 12, 42, 58, 52, 44, 33] can synthesize realistic human motions for HOIs for static objects only. They usually synthesize the motion in the last mile of interaction, i.e. the motion between the given starting human pose and the final interaction pose, and overlook the movement of the objects when the human is interacting with them. On the other hand, existing methods for motion generation with dynamic objects do not adequately reflect real-world complexity. For instance, they focus on grasping small objects [10], provide the object motion as conditioning [25], predict deterministic interactions between the human and the same object without the diversity [54, 37], consider only a small set of interactions (e.g. sit/lift [23], sit/lie down [12], sit [17, 58, 33], grasp [52, 44]), or investigate a single type of object (e.g. chair [17, 58]).
@@
In this paper, we go beyond and propose HOI-Diff for 3D HOIs synthesis involving humans interacting with different types of objects in diverse ways, which are both physically plausible and semantically faithful to the textual prompt, as shown in Fig. 1. Our key insight is to decompose 3D HOIs synthesis into three modules: (a) coarse 3D HOIs generation that extends the human motion diffusion model [45] to a dual-branch diffusion model (HOI-DM) to generate both human and object motions conditioning on the input text prompt. To encourage coherent motions, we develop a cross-attention communication module, exchanging information between the human and object motion generation models; (b) affordance prediction diffusion model (APDM) that estimates the contacting points between the human and object during the interactions driven by the textual prompt. Our APDM does not rely on the results of the HOI-DM and thus can recover from its potential errors. Moreover, it stochastically generates the contacting points to diversity the generated motions; and (c) affordance-guided interaction correction that incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects. Compared with designing a monolithic model, our proposed HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of the three modules.
@@
To our best knowledge, HOI-Diff is one of the first text-driven approach s capable of generating realistic motions for various HOIs and different types of objects during the interactions. While our approach pioneers this capability, we note the existence of concurrent works such as  [9, 24, 49], which also explore similar directions. For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [3] with text descriptions, which mitigates the issue of severe data scarcity for text-driven 3D HOIs generation. In addition, we evaluate our approach on the OMOMO dataset [25], which focuses on the manipulation of two hands. Extensive experiments validate the effectiveness and design choices of our approach, particularly for dynamic objects, thereby enabling a set of new applications in human motion generation.
@@ 2 Related Work
Human Motion Generation with Diffusion Models. The denoising diffusion models have been widely used 2D image generations [39, 40, 36] and achieved impressive results. Recent work [56, 45, 5, 18, 38, 1, 2, 4, 6, 41, 43, 46, 51, 55, 57, 59, 53] apply the diffusion model in the task of human motion generation. While these methods have successfully generated human motion, they usually generate isolated motions in the free space without considering the objects the human is interacting with. Our method is primarily focused on motion generation with human-object interactions.
@@
Scene- and Object-Aware Human Motion Generation. Recent works condition motion synthesis on scene geometry [15, 60, 48, 50]. This facilitates the understanding of human-scene interactions. However, the motion fidelity is compromised due to the lack of paired full scene-motion data. Other approaches [23, 17, 12, 42, 58, 33] instead focus on the interactions with the objects and can produce realistic motions. However, they focus on interacting with static objects with limited interactions. OMOMO [25] proposes a conditional diffusion framework that can generate full-body motion from the object motion. The object motion is needed as input in OMOMO, whereas our method can jointly synthesize human motion and object motion. IMoS [10] synthesizes the full-body human along with the 3D object motions from textual inputs. Although they can generate realistic interactions, IMoS only focuses on grasping small objects with hands. InterDiff [54] predicts whole-body interactions with dynamic objects. However, it focuses on the motion prediction task, where the future motion needs to be forecasted based on past motion information and object information. Note that the interaction type is deterministic, given the object and past motion information. Different from this, we tackle the motion synthesis task, where the interaction with the same object can be controlled by the text prompt. Recently, there has been a surge of interest in the text-driven synthesis of 3D human-object interactions for dynamic objects, resulting in the development of concurrent works [9, 24, 49].
Multi-Human Motion Generation. Recent works [41, 26, 41] incorporate human-to-human interactions into the motion diffusion process. They usually also adopt cross-attention to exchange information from different human motions. Different from them, our goal is to generate coarse HOIs instead of doing so in a single run, which alleviates the burden of the model.
@@
Affordance Estimation. The affordance estimation on 3D point cloud is studied in [30, 7, 22, 16, 29, 20, 21].  [20] formulates this task as the point cloud segmentation task from the extracted geometric features.  [21] voxelizes the point cloud and creates affordance maps with interactive manipulation.  [22, 16] focuses on the object grasping affordances and some specific applications.  [29] learns the affordance for object-object interactions.  [30] addressed the task of open-vocabulary affordance learning. Overall affordance learning is a very challenging task even the shape of point cloud is given. Instead of predicting the point-wise contact labels, in our method, we propose to make some hypotheses to simplify human-object interactions, making it more tractable without significantly compromising accuracy.
@@ 3 Method
3.1 Overview
Motion Representations. We denote a 3D HOI sequence as ${\bm{x}}=\{{\bm{x}}^{h},{\bm{x}}^{o}\}$. It consists of human motion sequence ${\bm{x}}^{h}\in\mathbb{R}^{L\times D^{h}}$ and object motion sequence ${\bm{x}}^{o}\in\mathbb{R}^{L\times D^{o}}$, where $L$ denotes the length of the sequence. For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [11] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels. For the object motion sequence ${\bm{x}}^{o}$, we assume the object geometry is given as an input, and thus we only need to estimate its 6DoF poses in the generation, i.e., $D^{o}=6$. We represent each object instance as a point cloud of 512 points ${\bm{p}}\in\mathbb{R}^{512\times 3}$.
@@ 3.1 Overview
where $t$ is the denoising step. Following [45], instead of predicting the noise $\mathbf{\epsilon}_{t}$ in each diffusion step, our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$.
Figure 2: Overview of HOI-Diff for 3D HOIs generation using diffusion models. Our key insight is to decompose the generation task into three modules: (a) coarse 3D HOI generation using a dual-branch diffusion model (HOI-DM), (b) affordance prediction diffusion model (APDM) to estimate the contacting points of humans and objects, and (c) afforance-guided interaction correction, which incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects to form coherent HOIs.
We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [31]
@@
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta_{t}\in(0,1)$ are the variance schedule. $\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [14] is a variance scheduler of choice. Similar to ${\bm{x}}_{t}$, ${\bm{\mu}}_{t}$ consists of ${\bm{\mu}}^{h}_{t}$ and ${\bm{\mu}}^{o}_{t}$, corresponding to human and object motion, respectively.
Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation. It has two branches designed for generating human and object motions individually. A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions. The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
Figure 3: Illustration of HOI-DM architecture for coarse 3D HOIs generation. It has two branches designed for generating human and object motions individually. A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions. The human motion model $M^{h}$ finetunes a pretrained human MDM [45].
@@
Fig. 2 shows the overview of our proposed approach. We introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM), which can produce diverse yet consistent motions, capturing the intricate interplay and mutual interactions between humans and objects (Sec. 3.2). To ensure physically plausible contact between humans and objects, we propose a novel affordance prediction diffusion model (APDM) (Sec. 3.3), whose output will be used as classifier guidance (Sec. 3.4) to correct the interactions at each diffusion step of human/object motion generation.
3.2 Coarse 3D HOIs Generation
First, we introduce a dual-branch Human-Object Interaction Diffusion Model (HOI-DM) to generate human and object motions that are roughly coherent. As shown in Fig. 4, it consists of two Transformer models [47], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [45]. Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predict clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
To enhance the learning of interactions of the human and object when generating their motion, we introduce a Communication Module ($CM$) designed for exchanging feature representations between the human MDM $M^{h}$ and the object MDM $M^{o}$. $CM$ is a Transformer block that receives the intermediate feature ${\bm{f}}^{h},\bm{f}^{o}$ from both $M^{h}$ and $M^{o}$. It then processes these inputs to generate refined updates based on the cross attention mechanism [47]. The updated feature representations $\tilde{{\bm{f}}_{h}}$ and $\tilde{{\bm{f}}_{o}}$ of the human and object are then conditioned on each other, which are then fed into the subsequent layers of their respective branches to estimate clean human and object motion ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively. The $CM$ is inserted at the 4th transformer layer for human MDM and the last layer for object MDM, which was empirically found to work better.
Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [45]. This fine-tuning is critical to ensure the smoothness of the generated human motions. We ablate this design choice in Sec. 4.3. Object MDM is trained from scratch. We modify the input and output linear layers to take in the object motion which has a different dimension from the human motion. More methodology details of HOI-DM can be found in the supplementary material.

### v3

# HOI-Diff: Text-Driven Synthesis of 3D Human-Object Interactions using Diffusion Models

@@ Abstract
We address the problem of generating realistic 3D human-object interactions (HOIs) driven by textual prompts. To this end, we take a modular design and decompose the complex task into simpler sub-tasks. We first develop a dual-branch diffusion model (DBDM) to generate both human and object motions conditioned on the input text, and encourage coherent motions by a cross-attention communication module between the human and object motion generation branches. We also develop an affordance prediction diffusion model (APDM) to predict the contacting area between the human and object during the interactions driven by the textual prompt. The APDM is independent of the results by the DBDM and thus can correct potential errors by the latter. Moreover, it stochastically generates the contacting points to diversify the generated motions. Finally, we incorporate the estimated contacting points into the classifier-guidance to achieve accurate and close contact between humans and objects. To train and evaluate our approach, we annotate the BEHAVE dataset with text descriptions. Experimental results on BEHAVE and OMOMO demonstrate that our approach produces realistic HOIs with various interactions and different types of objects.
Figure 1: HOI-Diff can generate realistic motions for 3D human-object interactions given a text prompt and object geometry. Please see the supplementary material for video results. Darker color indicates later frames in the sequence. Best viewed in color.
The first two authors contributed equally.
@@ 1 Introduction
The generation of natural and physically plausible 3D HOIs involves humans interacting with dynamic objects in various ways according to the text prompts, thereby posing several challenges. First, the variability of object shapes makes it particularly challenging to generate semantically meaningful contact between the human and object to avoid floating objects. Second, the generated HOIs should be faithful to the input text prompts as there are many plausible interactions between human and the same object (e.g, a person carries a chair, sits on a chair, pushes or pulls a chair). Text-driven 3D HOI synthesis with a diverse set of interactions is not yet fully addressed. Third, the development and evaluation of 3D HOI synthesis models requires a high-quality human motion dataset with various HOIs and textual descriptions, but existing datasets lack either diverse HOIs [Guo_2022_CVPR, kit2016, li2023controllable] or detailed textual descriptions with interacting body parts and action [bhatnagar22behave, diller2023cghoi]. It is important to note that CG-HOI [diller2023cghoi] has not made their code or annotations publicly available. In contrast, we will release both our code and annotations.
@@
Current methods cannot fully handle all the challenges. On one hand, recent methods [kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, wu2022saga, taheri2021goal, Pi_2023_ICCV] can synthesize realistic human motions for HOIs for static objects only. They usually synthesize the motion in the last mile of interaction, i.e, the motion between the given starting human pose and the final interaction pose, and overlook the movement of the objects when the human is interacting with them. On the other hand, existing methods for motion generation with dynamic objects do not adequately reflect real-world complexity. For instance, they focus on grasping small objects [ghosh2023imos], provide the object motion as conditioning [li2023object], predict deterministic interactions between the human and the same object without the diversity [xu2023interdiff, razali2023action], consider only a small set of interactions (e.g., sit/lift [kulkarni2023nifty], sit/lie down [hassan_samp_2021], sit [jiang2022chairs, zhang2022couch, Pi_2023_ICCV], grasp [wu2022saga, taheri2021goal]), or investigate a single type of object (e.g., chair [jiang2022chairs, zhang2022couch]).
@@
In this paper, we introduce HOI-Diff for 3D HOIs synthesis involving humans interacting with different types of objects in diverse ways, which are both physically plausible and semantically faithful to the textual prompt, as shown in Figure 1. Our key insight is to decompose 3D HOIs synthesis into three modules to reduce the complexity of this challenging task. (a) coarse 3D HOIs generation that extends the human motion diffusion model [tevet2023human] to a dual-branch diffusion model (DBDM) to generate both human and object motions conditioning on the input text prompt. To encourage coherent motions, we develop a cross-attention communication module, exchanging information between the human and object motion generation models; (b) affordance prediction diffusion model (APDM) that estimates the contacting points between the human and object during the interactions driven by the textual prompt. Our APDM does not rely on the results of the DBDM and thus can recover from its potential errors. Moreover, it stochastically generates the contacting points to diversity the generated motions; and (c) affordance-guided interaction correction that incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects, significantly alleviating the cases of floating objects. Compared with designing a monolithic model, HOI-Diff disentangles motion generation for humans and objects and estimation of their contacting points, which are later integrated to form coherent and diverse HOIs, reducing the complexity and burden for each of the three modules.
@@
For both training and evaluation purposes, we annotate each video sequence in BEHAVE dataset [bhatnagar22behave] with text descriptions, which mitigates the issue of severe data scarcity for text-driven 3D HOIs generation. In addition, we evaluate our approach on the OMOMO dataset [li2023object], which focuses on the manipulation of two hands. Extensive experiments validate the effectiveness and design choices of our approach, particularly for dynamic objects, thereby enabling a set of new applications in human motion generation.
@@ 2 Related Work
Human Motion Generation with Diffusion Models. The denoising diffusion models have been widely used 2D image generations [rombach2022high, saharia2022photorealistic, ramesh2021zero] and achieved impressive results. Recent work [zhang2022motiondiffuse, tevet2023human, chen2023executing, karunratanakul2023gmd, rempeluo2023tracepace, ahn2023can, barquero2023belfusion, chen2023humanmac, dabral2023mofusion, shafir2023human, sun2023towards, tian2023transfusion, wei2023understanding, zhang2023generating, zhang2023remodiffuse, zhang2023tedi, xie2023omnicontrol] apply the diffusion model in the task of human motion generation. While these methods have successfully generated human motion, they usually generate isolated motions in the free space without considering the objects the human is interacting with. Our method is primarily focused on motion generation with human-object interactions.
@@
Scene- and Object-Aware Human Motion Generation. Recent works condition motion synthesis on scene geometry [huang2023diffusion, zhao2023synthesizing, wang2022towards, wang2022humanise]. This facilitates the understanding of human-scene interactions. However, the motion fidelity is compromised due to the lack of paired full scene-motion data. Other approaches p[kulkarni2023nifty, jiang2022chairs, hassan_samp_2021, starke2019neural, zhang2022couch, Pi_2023_ICCV] instead focus on the interactions with the objects and can produce realistic motions. However, they focus on interacting with static objects with limited interactions. OMOMO [li2023object] can generate full-body motion from the object motion. The object motion is needed as input in OMOMO, whereas our method can jointly synthesize human motion and object motion. IMoS [ghosh2023imos] synthesizes the full-body human along with the 3D object motions from textual inputs, but it only focuses on grasping small objects with hands. InterDiff [xu2023interdiff] predicts whole-body interactions with dynamic objects. Note that the interaction type is deterministic. Different from this, we tackle the motion synthesis task, where the interaction with the same object can be controlled by the text prompt. Recently, there has been a surge of interest in the text-driven synthesis of 3D human-object interactions for dynamic objects, resulting in the development of concurrent works [diller2023cghoi, wang2023physhoi, li2023controllable, song2024hoianimator, xu2024interdreamer]. CG-HOI [diller2023cghoi] and HOIAnimator [song2024hoianimator] uses SMPL parameters as the motion representation, which may result in unsmooth motion due to the potential difficulty in optimization. Instead, we use common skeletal joints similar to most text-to-motion methods, harnessing the power of pre-trained human motion generation models. Chois li2023controllable relies on the initial state and object waypoints to generate HOIs, which reduces motion diversity for both the human and the object. InterFusion [dai2024interfusion] and F-HOI [yang2024f] generate static 3D HOIs from text description, lacking both human and object motions.
Figure 2: Overview of HOI-Diff for 3D HOIs generation using diffusion models. Our key insight is to decompose the generation task into three modules: (a) coarse 3D HOI generation using a dual-branch diffusion model (DBDM), (b) affordance prediction diffusion model (APDM) to estimate the contacting points of humans and objects, and (c) affordance-guided interaction correction, which incorporates the estimated contacting information and employs the classifier-guidance to achieve accurate and close contact between humans and objects to form coherent HOIs.
@@
Affordance Estimation. The affordance estimation on 3D point cloud is studied in ngyen2023open, deng20213d, kokic2017affordance, iriondo2021affordance, mo2022o2o, kim2014semantic, kim2015interactive. Overall affordance learning is a very challenging task. Instead of predicting the point-wise contact labels, we simplify it by directly regressing the contact points for human-object interactions, making it more tractable without significantly compromising accuracy.
@@ 3 Method
The overview of our proposed approach are illustrated in Figure 2. We introduce a dual-branch Human-Object Interaction Diffusion Model (DBDM), which can produce diverse yet consistent motions, capturing the intricate interplay and mutual interactions between humans and objects (Sec. 3.2). To ensure physically plausible contact between humans and objects, we propose a novel affordance prediction diffusion model (APDM) (Sec. 3.3), whose output will be used as classifier guidance (Sec. 3.4) to correct the interactions at each diffusion step of human/object motion generation.
3.1 Background
Motion Representations. We denote a 3D HOI sequence as ${\bm{x}}=\{{\bm{x}}^{h},{\bm{x}}^{o}\}$. It consists of human motion sequence ${\bm{x}}^{h}\in\mathbb{R}^{L\times D^{h}}$ and object motion sequence ${\bm{x}}^{o}\in\mathbb{R}^{L\times D^{o}}$, where $L$ denotes the length of the sequence. For ${\bm{x}}^{h}$, we adopt the redundant representation widely used in human motion generation [Guo_2022_CVPR] with $D^{h}=263$, which include pelvis velocity, local joint positions, velocities and rotations of other joints in the pelvis space, and binary foot-ground contact labels. For the object motion sequence ${\bm{x}}^{o}$, we assume the object geometry is given as an input, and thus we only need to estimate its 6DoF poses in the generation, i.e, $D^{o}=6$. We represent each object instance as a point cloud of 512 points ${\bm{p}}\in\mathbb{R}^{512\times 3}$.
@@ 3.1 Background
where $t$ is the denoising step. Following [tevet2023human], our diffusion model $M_{\theta}$ with parameters $\theta$ predicts the final clean motion ${\bm{x}}_{0}=M_{\theta}({{\bm{x}}_{t}},t,{\bm{c}})$.
We sample $\mathbf{x}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{t},\Sigma_{t})$ and compute the mean as in [nichol2021improved]
@@
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta_{t}\in(0,1)$ are the variance schedule. $\Sigma_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t}$ [Ho2020-ew] is a variance scheduler of choice. Similar to ${\bm{x}}_{t}$, ${\bm{\mu}}_{t}$ consists of ${\bm{\mu}}^{h}_{t}$ and ${\bm{\mu}}^{o}_{t}$, corresponding to human and object motion, respectively.
3.2 Coarse 3D HOIs Generation
First, we introduce a dual-branch diffusion model (DBDM) to generate human and object motions that are roughly coherent. As shown in  Figure 4, it consists of two Transformer models [vaswani2017attention], human motion diffusion model (MDM) $M^{h}$ and object MDM $M^{o}$, which work similar to [tevet2023human]. Specifically, at the diffusion step $t$, they take the text description and noisy motions ${\bm{x}}_{t}^{h}$ and ${\bm{x}}_{t}^{o}$ as input and predict clean human and object motions ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively.
To enhance the learning of interactions of the human and object when generating their motion, we introduce a Communication Module ($CM$) designed for exchanging feature representations between the human MDM $M^{h}$ and the object MDM $M^{o}$. $CM$ is a Transformer block that receives the intermediate feature ${\bm{f}}^{h},\bm{f}^{o}$ from both $M^{h}$ and $M^{o}$. It then processes these inputs to generate refined updates based on the cross attention mechanism [vaswani2017attention]. The updated feature representations $\tilde{{\bm{f}}_{h}}$ and $\tilde{{\bm{f}}_{o}}$ of the human and object are then conditioned on each other, which are then fed into the subsequent layers of their respective branches to estimate clean human and object motion ${\bm{x}}_{0}^{h}$ and ${\bm{x}}_{0}^{o}$, respectively. The $CM$ is inserted at the 4th transformer layer for human MDM and the last layer for object MDM, which was empirically found to work better.
Given the limited data availability for 3D HOI generation, during training, the human motion model $M^{h}$ finetunes a pretrained human MDM [tevet2023human]. This fine-tuning is critical to ensure the smoothness of the generated human motions. We ablate this design choice in Sec. LABEL:ablation. Object MDM is trained from scratch. We modify the input and output linear layers to take in the object motion which has a different dimension from the human motion. More details of DBDM are in Appendix A.1.
[…]
Figure 3: Illustration of DBDM architecture for coarse 3D HOIs generation. It has two branches designed for generating human and object motions individually. A mutual cross-attention is introduced to allow information exchange between two branches to generate coherent motions. The human motion model $M^{h}$ finetunes a pretrained MDM [tevet2023human].
[…]